OpenStudy (anonymous):

if (3/4 -2/3)+(1/2+1/3) is calculated and the answer reduced to simplest terms what is the denominator of the resulting fraction

OpenStudy (radar):

Remove the parenthesis giving 3/4-2/3+1/2+1/3 You can combine those fractions that have a common denominator the -2/3 and the 1/3 becomes -1/3 Now the combining is of 3/4-1/3+1/2 find the LCD for 4, 3, and 2. This can be solved by multiplying the 4 and 3 (as 2 is already contained in the 4) Now the fractions are 9/12-4/12+6/12=15/12-4/12=11/12, the denominator is 12 and 11/12 is in its simplest terms.

OpenStudy (radar):

Are there any question or concerns?